# DMUS-MP-25/04

### v1

@@ 2.1 Relation with massless pure-RR $AdS_{3}\times S^{3}\times T^{4}$
The Sine-Gordon $S$-matrix at the special value of the coupling discussed in the previous section reduces to a known $S$-matrix discussed by Fendley and Intriligator [44, 45, 46]. This $S$-matrix is almost the same as the massless [^5] For definiteness, we shall always mean left-left same-chirality massless scattering for the sake of this paper. $AdS_{3}\times S^{3}\times T^{4}$, but not quite. It turns out that it differs from it by some factors of $\pm i$ [14], and this is because it has the exact same ${\cal{N}}=2$ supersymmetry which we have just described, but the fermionic assignement of the states is different: in fact in $AdS_{3}$ the fermionic number is the more traditional $|\phi|=0$, and $|\psi|=-1$ (or, equivalently, $|\psi|=1$ mod 22) [^6] This fermionic number is the one used in [53], which therefore ends up being exactly like massless $AdS_{3}$ although apparently for the wrong reason (see comment at the very end of section 3 of [45]).. Let us summarise here below the situation - notice that we use everywhere in this paper the symbol $\theta_{i}$ for the particle’s relativistic rapidity, even if in the case of massless pure RR $AdS_{3}$ we mean the relativistic $\gamma_{i}$ variable [34]. Without any index, $\theta$ always indicates the rapidity-difference.
@@
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&-i\tanh\frac{\theta}{2}&0\\ 0&-i\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (2.11)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{E},\quad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{F},\qquad|\phi|=\frac{1}{2},\quad|\psi|=-\frac{1}{2},\quad|\hat{E}|=1,\quad|\hat{F}|=-1,$
@@
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&\tanh\frac{\theta}{2}&0\\ 0&-\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (2.12)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{E},\quad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{F},\quad|\phi|=0,\qquad|\psi|=-1,\quad|\hat{E}|=1,\quad|\hat{F}|=-1,$
@@
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&-i\tanh\frac{\theta}{2}&0\\ 0&-i\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (2.13)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{-2H}\otimes\hat{E},\quad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{2H}\otimes\hat{F},\quad|\phi|=|\psi|=0,\quad|\hat{E}|=|\hat{F}|=0,\quad q_{0}=-i.$
@@
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&\tanh\frac{\theta}{2}&0\\ 0&-\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (2.14)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}\otimes\hat{E},\quad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}\otimes\hat{F}$,
    $|\phi|=|\psi|=0,\quad|\hat{E}|=|\hat{F}|=0,\quad q_{0}=-i.$
@@
$\Delta\hat{E}=\hat{E}\otimes\mathbbmss{1}+{q_{0}}^{2H-\mathbbmss{1}}\otimes\hat{E}$. (2.15)
@@
$\Delta\hat{E}|\alpha\rangle\otimes|\psi\rangle=(\hat{E}\otimes\mathbbmss{1}+q^{2H-\mathbbmss{1}}\otimes\hat{E})|\phi\rangle\otimes|\psi\rangle=e|\alpha\rangle\otimes|\psi\rangle+(-1)^{\sigma_{3}-\mathbbmss{1}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle,\qquad\alpha=\phi,\psi.$
Both $|\phi\rangle=\begin{pmatrix}1\\ 0\end{pmatrix}$ and $|\psi\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$ have fermionic number $|\phi|=|\psi|=0$, and $\hat{E}$ has zero fermionic number - because it sends the state $|\psi\rangle$ into the state $|\phi\rangle$. We can now rephrase (2.7) in this way:
$\Delta(\hat{E})|\phi\rangle\otimes|\psi\rangle=(\mathbbmss{1}\otimes\hat{E}+\hat{E}\otimes\mathbbmss{1})|\phi\rangle\otimes|\psi\rangle=x_{2}\,|\phi\rangle\otimes|\phi\rangle+\hat{E}|\phi\rangle\otimes|\psi\rangle,$
$\Delta(\hat{E})|\psi\rangle\otimes|\psi\rangle=(\mathbbmss{1}\otimes\hat{E}+\hat{E}\otimes\mathbbmss{1})|\psi\rangle\otimes|\psi\rangle=-x_{2}\,|\psi\rangle\otimes|\phi\rangle+\hat{E}|\psi\rangle\otimes|\psi\rangle.$
First, $\hat{E}$ is represented as $x_{i}E_{\phi\psi}=e^{\frac{\theta_{i}}{2}}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$, and has fermionic number $|\hat{E}|=1$; second, $|\phi\rangle$ is represented as $\begin{pmatrix}1\\ 0\end{pmatrix}$, and has fermionic number $|\phi|=0$; third, $|\psi\rangle$ is represented as $\begin{pmatrix}0\\ 1\end{pmatrix}$, and has fermionic number $|\psi|=-1$. The other cases work out analogously. Basically this time the quantum-group deformation $q^{2H-\mathbbmss{1}}$, resp. $q^{-2H+\mathbbmss{1}}$, in the coproduct at this value of the coupling become equal to $(-1)^{|\hat{E}|\begin{pmatrix}0&0\\ 0&-1\end{pmatrix}}$, resp. $(-1)^{|\hat{F}|\begin{pmatrix}0&0\\ 0&-1\end{pmatrix}}$, which simulates ordinary supersymmetry. We have summarised the situation in fig. 1.
Figure 1: Setting the coupling to a special value in Sine-Gordon produces the Fendley-Intriligator model, which has the same symmetry as massless $AdS_{3}$ but different assignment of fermionic numbers of the states.
2.1.1 Going fractional
It is convenient to obtain an $R$-matrix which depends on an unspecified fractional fermionic grading in such a way to interpolate between Sine-Gordon and $AdS_{3}$ when q=iq=-i:
@@ 2.1.1 Going fractional
$S_{a}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{\pi\theta}{\xi}}(q^{2}-1)}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&\frac{-iqe^{-i\pi\alpha}\big{(}e^{\frac{2\pi\theta}{\xi}}-1\big{)}}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&0\\ 0&\frac{iqe^{i\pi\alpha}\big{(}e^{\frac{2\pi\theta}{\xi}}-1\big{)}}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&\frac{e^{\frac{\pi\theta}{\xi}}(q^{2}-1)}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&0\\ 0&0&0&1\end{pmatrix}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&S_{R}&ie^{-i\pi\alpha}S_{T}&0\\ 0&-ie^{i\pi\alpha}S_{T}&S_{R}&0\\ 0&0&0&1\end{pmatrix}$,
$\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+e^{-2H(\log[q]-\log[-i])+\mbox{fracto}}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+e^{2H(\log[q]-\log[-i])-\mbox{fracto}}\otimes\hat{F}$,
$|\phi|=0,\qquad|\psi|=0,\quad|\hat{E}|=0,\quad|\hat{F}|=0,\qquad\mbox{fracto}=\begin{pmatrix}\alpha&0\\ 0&\alpha-1\end{pmatrix}$,
with $\hat{E}=e^{\frac{\pi\theta_{i}}{\xi}}E_{\phi\psi}$, $\hat{F}=e^{\frac{\pi\theta_{i}}{\xi}}E_{\psi\phi}$, and where $S_{R}$, $S_{T}$ are the Sine-Gordon entries (2.2). One can verify that
@@
We notice that $S_{\alpha}$ solves the Yang-Baxter equation (YBE) for all $\alpha$, and is braiding-unitary $S_{\alpha}(\theta)S_{\alpha}(-\theta)=\mathbbmss{1}\otimes\mathbbmss{1}$ if $\Phi(\theta)\Phi(-\theta)=1$. The parameter $(\alpha,\alpha-1)$ is effectively the statistics of particles $(\phi,\psi)$, which we are using to interpolate between Sine-Gordon (Fendley-Intriligator) $(\frac{1}{2},-\frac{1}{2})$ and $AdS_{3}$ (0,1)(0,-1), as $\alpha$ goes from $\frac{1}{2}$ to 0. We have summarised the situation in fig. 2. $S_{\alpha}$ is also physical-unitary, i.e. $S_{\alpha}$ is a unitary matrix for real rapidities, if $\Phi(\theta)$ is a pure phase for real $\theta$ (and $\alpha$ is real.)
Figure 2: The fractional interpolation connects the diagram 1 which previously was disconnected.
3 Relaxing the special coupling

### v4

@@ II.1 Relation with massless pure-RR $AdS_{3}\times S^{3}\times T^{4}$
The Sine-Gordon $S$-matrix at the special value of the coupling discussed in the previous section reduces to a known $S$-matrix discussed by Fendley and Intriligator [45, 46, 47]. This $S$-matrix is almost the same as the massless  [^6] For definiteness, we focus on right-moving kinematics. $AdS_{3}\times S^{3}\times T^{4}$, but not quite. It turns out that it differs from it by some factors of $\pm i$ [14], and this is because it has the exact same ${\cal{N}}=2$ supersymmetry which we have just described, but the fermionic assignement of the states is different: in fact in $AdS_{3}$ the fermionic number is the more traditional $|\phi|=0$, and $|\psi|=-1$ (or, equivalently, $|\psi|=1$ mod $2$)  [^7] This fermionic number is the one used in [56], which therefore ends up being exactly like massless $AdS_{3}$ although for an incorrect reason (see comment at the very end of section 3 of [46]).. Let us summarise here below the situation - notice that we use everywhere in this paper the symbol $\theta_{i}$ for the particle’s relativistic rapidity, even if in the case of massless pure RR $AdS_{3}$ we mean the relativistic $\gamma_{i}$ variable [35]. Without any index, $\theta$ always indicates the rapidity-difference.
@@
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&-i\tanh\frac{\theta}{2}&0\\ 0&-i\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$, (12)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{F}$,
    $\hskip 18.49988pt|\phi|=\frac{1}{2},\qquad|\psi|=-\frac{1}{2},\qquad|\hat{E}|=1,\qquad|\hat{F}|=-1,$
@@
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&\tanh\frac{\theta}{2}&0\\ 0&-\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$, (13)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{F}$,
    $\qquad|\phi|=0,\hskip 18.49988pt|\psi|=-1,\qquad|\hat{E}|=1,\qquad|\hat{F}|=-1,$
@@
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&-i\tanh\frac{\theta}{2}&0\\ 0&-i\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$, (14)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{-2H}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{2H}\otimes\hat{F}$,
    $\qquad|\phi|=|\psi|=0,\qquad|\hat{E}|=|\hat{F}|=0,\qquad q_{0}=-i.$
@@
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&\tanh\frac{\theta}{2}&0\\ 0&-\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$, (15)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}\otimes\hat{F}$,
    $|\phi|=|\psi|=0,\qquad|\hat{E}|=|\hat{F}|=0,\qquad q_{0}=-i.$
@@
$\Delta\hat{E}=\hat{E}\otimes\mathbbmss{1}+{q_{0}}^{2H-\mathbbmss{1}}\otimes\hat{E}$. (16)
@@
$\Delta\hat{E}|\alpha\rangle\otimes|\psi\rangle=(\hat{E}\otimes\mathbbmss{1}+q^{2H-\mathbbmss{1}}\otimes\hat{E})|\alpha\rangle\otimes|\psi\rangle=$
$\hat{E}|\alpha\rangle\otimes|\psi\rangle+(-1)^{\sigma_{3}-\mathbbmss{1}}\,x_{2}\,|\alpha\rangle\otimes|\phi\rangle,\hskip 18.49988pt\alpha=\phi,\psi.$
Both $|\phi\rangle=\begin{pmatrix}1\\ 0\end{pmatrix}$ and $|\psi\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$ have fermionic number $|\phi|=|\psi|=0$, and $\hat{E}$ has zero fermionic number - because it sends the state $|\psi\rangle$ into the state $|\phi\rangle$. We can now rephrase (8) in this way:
$\Delta(\hat{E})|\phi\rangle\otimes|\psi\rangle=(\mathbbmss{1}\otimes\hat{E}+\hat{E}\otimes\mathbbmss{1})|\phi\rangle\otimes|\psi\rangle$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt=x_{2}\,|\phi\rangle\otimes|\phi\rangle+\hat{E}|\phi\rangle\otimes|\psi\rangle,$
$\Delta(\hat{E})|\psi\rangle\otimes|\psi\rangle=(\mathbbmss{1}\otimes\hat{E}+\hat{E}\otimes\mathbbmss{1})|\psi\rangle\otimes|\psi\rangle$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt=-x_{2}\,|\psi\rangle\otimes|\phi\rangle+\hat{E}|\psi\rangle\otimes|\psi\rangle.$
First, $\hat{E}$ is represented as $x_{i}E_{\phi\psi}=e^{\frac{\theta_{i}}{2}}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$, and has fermionic number $|\hat{E}|=1$; second, $|\phi\rangle$ is represented as $\begin{pmatrix}1\\ 0\end{pmatrix}$, and has fermionic number $|\phi|=0$; third, $|\psi\rangle$ is represented as $\begin{pmatrix}0\\ 1\end{pmatrix}$, and has fermionic number $|\psi|=-1$. The other cases work out analogously. Basically this time the quantum-group deformation $q^{2H-\mathbbmss{1}}$, resp. $q^{-2H+\mathbbmss{1}}$, in the coproduct at this value of the coupling become equal to $(-1)^{|\hat{E}|\Sigma}$, resp. $(-1)^{|\hat{F}|\Sigma}$, with $\Sigma=\begin{pmatrix}0&0\\ 0&-1\end{pmatrix}$, which simulates ordinary supersymmetry.
II.1.1 Going fractional
It is convenient to obtain an $S$-matrix which depends on an unspecified fractional fermionic grading in such a way to interpolate between Sine-Gordon and $AdS_{3}$ when q=iq=-i:
@@ II.1.1 Going fractional
$S_{\alpha}=\Phi_{\alpha}(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{\pi\theta}{\xi}}(q^{2}-1)}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&\frac{-iqe^{-i\pi\alpha}\big(e^{\frac{2\pi\theta}{\xi}}-1\big)}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&0\\ 0&\frac{iqe^{i\pi\alpha}\big(e^{\frac{2\pi\theta}{\xi}}-1\big)}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&\frac{e^{\frac{\pi\theta}{\xi}}(q^{2}-1)}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&0\\ 0&0&0&1\end{pmatrix}$=
$\Phi_{\alpha}(\theta)\begin{pmatrix}1&0&0&0\\ 0&S_{R}&ie^{-i\pi\alpha}S_{T}&0\\ 0&-ie^{i\pi\alpha}S_{T}&S_{R}&0\\ 0&0&0&1\end{pmatrix}$,
$\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+e^{-2H(\log[q]-\log[-i])+\mbox{fracto}}\otimes\hat{E}$,
$\hskip 18.49988pt\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+e^{2H(\log[q]-\log[-i])-\mbox{fracto}}\otimes\hat{F}$,
$|\phi|,|\psi|=0,\qquad|\hat{E}|=0,\qquad|\hat{F}|=0,$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\mbox{fracto}=\log(-1)\begin{pmatrix}\alpha&0\\ 0&\alpha-1\end{pmatrix}$, (17)
with $\hat{E}=e^{\frac{\pi\theta_{i}}{\xi}}E_{\phi\psi}$, $\hat{F}=e^{\frac{\pi\theta_{i}}{\xi}}E_{\psi\phi}$, and where $S_{R}$, $S_{T}$ are the Sine-Gordon entries (3). One can verify that
@@
We notice that $S_{\alpha}$ solves the Yang-Baxter equation (YBE) for all $\alpha$, and is braiding-unitary $S_{\alpha}(\theta)S_{\alpha}(-\theta)=\mathbbmss{1}\otimes\mathbbmss{1}$ if $\Phi_{\alpha}(\theta)\Phi_{\alpha}(-\theta)=1$. The parameter $(\alpha,\alpha-1)$ is effectively the statistics of particles $(\phi,\psi)$, which we are using to interpolate between Sine-Gordon (Fendley-Intriligator) $(\frac{1}{2},-\frac{1}{2})$ and $AdS_{3}$ (0,1)(0,-1), as $\alpha$ goes from $\frac{1}{2}$ to 0. $S_{\alpha}$ is also physical-unitary, i.e. $S_{\alpha}$ is a unitary matrix for real rapidities, if $\Phi(\theta)$ is a pure phase for real $\theta$ (and $\alpha$ is real.)
When $S_{\alpha}$ is taken for generic $q$ and $\alpha=\frac{1}{2}$, it reduces to the ordinary Sine-Gordon $S$-matrix (2). This means that $S_{\alpha}$ corresponds to a bona fide deformation of Sine-Gordon theory for particles with varying fermionic number - in the sense that one can dial the fermionic number as an independent real parameter. In this case we have also managed to derive the crossing-symmetry equation and solve it. To this purpose, as it is the case in $AdS_{3}$ literature, a second pair of states $(\bar{\phi},\bar{\psi})$ is introduced, and a new $S$-matrix scattering barred states $(\bar{\phi},\bar{\psi})$ against un-barred states $(\phi,\psi)$. We worked it out completely for q=iq=-i, the associated $R$-matrix being
$R_{\tiny\mbox{bar-unbar}}=\Phi_{mix}(\theta)\begin{pmatrix}-\tanh\frac{\theta}{2}&0&0&-\mbox{sech}\frac{\theta}{2}\\ 0&e^{-i\pi\alpha}&0&0\\ 0&0&-e^{i\pi\alpha}&0\\ -\mbox{sech}\frac{\theta}{2}&0&0&\tanh\frac{\theta}{2}\end{pmatrix}$,
such that the crossing equation is
$R_{\alpha}(\theta)(C^{-1}\otimes\mathbbmss{1})R^{\mbox{tr}_{1}}_{\tiny\mbox{bar-unbar}}(\theta+i\pi)(C\otimes\mathbbmss{1})=\mathbbmss{1}\otimes\mathbbmss{1},$
where $C=\begin{pmatrix}1&0\\ 0&i\end{pmatrix}$ is the charge-conjugation matrix in one of its possible versions, and $\mbox{tr}_{1}$ denotes transposition on the first space of the tensor product. Explicit evaluation of this crossing equation produces a condition for the dressing factors:
$\Phi_{\alpha}(\theta)\Phi_{mix}(\theta+i\pi)=-\tanh\frac{\theta}{2},$ (18)
which is independent on $\alpha$. This relation therefore admits as a possible solution the same formula as in one of the early treatments of pure massless $AdS_{3}$ [14], see also [21]:
$\Phi_{mix}(\theta)=i\Phi_{Z}(\theta),\hskip 18.49988pt\Phi_{\alpha}(\theta)=\Phi_{Z}(\theta)$ (19)
$\Phi_{Z}(\theta)=\prod_{\ell=1}^{\infty}\frac{\Gamma^{2}(\ell-\tau)\,\Gamma(\frac{1}{2}+\ell+\tau)\,\Gamma(-\frac{1}{2}+\ell+\tau)}{\Gamma^{2}(\ell+\tau)\,\Gamma(\frac{1}{2}+\ell-\tau)\,\Gamma(-\frac{1}{2}+\ell-\tau)},$ (20)
where
$\tau\equiv\frac{\theta}{2\pi i}.$ (21)
A complete analysis of poles and zeroes of this dressing factor, and the complete list of singularities of the $S$-matrix, is contained for example in [30]. In particular, no bound-state poles are present in the physical strip of the two-particle $S$-matrix, in accordance with the expectations for massless scattering.
III Relaxing the special coupling